Operations Report – February 14th

SOL: 4

Name of person filing report: Verónica Triviño

Non-nominal systems: None

Notes on non-nominal systems: N/A

Generator (hours run): Turned on at 1730 (14th Feb), turned off at 1500 (14th Feb)

Solar— SOC % – Turned off Gen (14th Feb) 83%, Turned on Gen (14th Feb) 67%

Diesel Reading – 97%

Propane Reading – ~ 50%

Ethanol Free Gasoline – 10 gallons

Water (Auxillary tank) – Not in use

Water (Static tank) – around 68%, ~ 372 gallons

Auxillary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab – ~ 252.5 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01412244 units

Toilet tank emptied: No

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Not Used

Hours: 50.1 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Not Used

Hours: 74.3 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Spirit rover used: Not Used

Hours: 73.7

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: N/A

ATV’s Used: N/A

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? Yes, going to town

Crew Car used and why, where? Not used

General notes and comments: Nothing to report

Summary of internet: Around noon, D. Masaitis (CO) checked the internet data status and saw a remaining data allocation of 392 MB. 20 minutes later, the data allocation dropped to 0 MB. Upon inspection of the crew’s devices, no offending device was found. A survey of the devices connected to the internet revealed only the router and D. Masaitis’ laptop, which was being used to correspond with Mission Control. The cause of the anomaly is still unknown and we will continue to monitor the Hab network.

Summary of suits and radios: Nothing to report.

Summary of Hab operations: Nothing to report

Summary of Green Hab operations: Nothing to report

Summary of Science Dome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Temperature in the upper deck of the Hab rose during the morning, at some points getting close to reaching 27 deg C, even though the heater in the lower deck was set by the engineer to slightly below the "ideal" middle setting of "3" (which usually gives the Hab a temperature of 21 – 23 deg C). The assistant director recommended turning the heater off and on again. After turning the heater off, the engineer and commander struggled to turn it back on again. As of writing this report, it is currently still off, however the temperature in the upper deck is a comfortable 23.8 deg C.